# UniCATS: A Unified Context-Aware Text-to-Speech Framework with Contextual VQ-Diffusion and Vocoding

### 20 Dec 2023 v4

@@ 2 UniCATS
Ground-truth - - 4.91 $\pm$ 0.04 4.51 $\pm$ 0.08 0.843
Encodec Acoustic token - 4.39 $\pm$ 0.07 4.00 $\pm$ 0.08 0.834
HifiGAN Semantic token X-vector 4.30 $\pm$ 0.08 3.96 $\pm$ 0.08 0.764
AudioLM Semantic token AR continuation 3.99 $\pm$ 0.07 3.96 $\pm$ 0.08 0.787
CTX-vec2wav Semantic token Contextual vocoding 4.75 $\pm$ 0.06 4.50 $\pm$ 0.07 0.840
@@ 3 Experiments and Results
Ground-truth 4.89 $\pm$ 0.04 4.54 $\pm$ 0.08 0.832 4.91 $\pm$ 0.04 4.50 $\pm$ 0.08 0.843
FastSpeech 2 3.81 $\pm$ 0.08 3.94 $\pm$ 0.06 0.820 3.65 $\pm$ 0.08 3.65 $\pm$ 0.07 0.765
VALL-E 4.23 $\pm$ 0.07 3.98 $\pm$ 0.06 0.795 4.17 $\pm$ 0.09 3.83 $\pm$ 0.07 0.775
UniCATS 4.54 $\pm$ 0.07 4.56 $\pm$ 0.07 0.831 4.43 $\pm$ 0.08 4.25 $\pm$ 0.08 0.832
@@ Appendix A Word Error Rate of Resynthesis
In this work, we opt to utilize vq-wav2vec tokens that retain richer prosody information than other semantic tokens such as HuBERT. In this section, we dive into evaluating the word error rate (WER) of resynthesis from HuBERT and vq-wav2vec. To this end, we train two CTX-vec2wav models with the two types of semantic tokens respectively. Then we resynthesize the speech in test set B from HuBERT and vq-wav2vec tokens and calculate their WERs respectively with Whisper (Radford et al. 2023), a well-known automatic speech recognition (ASR) model. The results are shown in the table below.

### 18 Jan 2024 v5

@@ 2 UniCATS
Ground-truth - - 4.91 $\pm$ 0.04 4.51 $\pm$ 0.08 0.837
Encodec Acoustic token - 4.39 $\pm$ 0.07 4.00 $\pm$ 0.08 0.829
HifiGAN Semantic token X-vector 4.30 $\pm$ 0.08 3.96 $\pm$ 0.08 0.776
AudioLM Semantic token AR continuation 3.99 $\pm$ 0.07 3.96 $\pm$ 0.08 0.801
CTX-vec2wav Semantic token Contextual vocoding 4.75 $\pm$ 0.06 4.50 $\pm$ 0.07 0.845
@@ 3 Experiments and Results
Ground-truth 4.89 $\pm$ 0.04 4.54 $\pm$ 0.08 0.833 4.91 $\pm$ 0.04 4.50 $\pm$ 0.08 0.837
FastSpeech 2 3.81 $\pm$ 0.08 3.94 $\pm$ 0.06 0.820 3.65 $\pm$ 0.08 3.65 $\pm$ 0.07 0.770
VALL-E 4.23 $\pm$ 0.07 3.98 $\pm$ 0.06 0.796 4.17 $\pm$ 0.09 3.83 $\pm$ 0.07 0.786
UniCATS 4.54 $\pm$ 0.07 4.56 $\pm$ 0.07 0.831 4.43 $\pm$ 0.08 4.25 $\pm$ 0.08 0.836
@@ Appendix A Word Error Rate of Resynthesis
In this work, we opt to utilize vq-wav2vec tokens that retain richer prosody information than other semantic tokens such as HuBERT. In this section, we dive into evaluating the word error rate (WER) of resynthesis from HuBERT and vq-wav2vec. To this end, we train two CTX-vec2wav models with the two types of semantic tokens respectively. Then we resynthesize the speech in test set B from HuBERT and vq-wav2vec tokens and calculate their WERs respectively with Whisper (Radford et al. 2023), a well-known automatic speech recognition (ASR) model. The results are shown in Table 4.